# BNSR-Invariants of Surface Houghton Groups

### v1

@@ 2.3 Characters and $\Sigma$-Invariants
[…]
For any group $G$, we define a character of $G$ to be a homomorphism from $G$ into the additive group $\mathbb{R}$. As characters factor through $G^{ab}$ (the abelianization of $G$), the space of characters is a vector space with the same dimension as the rank of $G^{ab}$. Excluding the trivial character and modding out by positive scaling, we obtain the Character Sphere $\Sigma(G)$. The $\Sigma$-invariants (also called BNS or BNSR invariants, for Bieri-Neumann-Strebel(-Renz)) are a filtration of the character sphere into subspaces
@@
Let $G$ be a group of type $F_{m}$, and let $Y$ be an $(m-1)$-connected CW complex on which $G$ acts cocompactly. Suppose that the stabilizer of any $k$-cell is of type $F_{m-k}$ and is contained in the kernel of every character of $G$ [^3] That the cell-stabilizers are contained in the kernels is the modification. For each non-trivial $\chi\in\operatorname{Hom}(G,\mathbb{R})$, there is a character height function $h_{\chi}:Y\rightarrow\mathbb{R}$, a continuous map satisfying $h_{\chi}(gy)=\chi(g)+h_{\chi}(y)$ for all $y\in Y$ and $g\in G$. Then $\Sigma^{m}(G)$ is the set of those $[\chi]$ such that the filtration $(Y^{t\leq h_{\chi}})_{t\in\mathbb{R}}$ is essentially $(m-1)$-connected.
Of what use are these invariants, one might ask? A standard result is the following classification of when finiteness properties of $G$ are preserved to subgroups containing the commutator.
@@
Let $G$ be a group of type $F_{m}$, and let $K$ be a subgroup so that $[G,G]\leq K\leq G$. Then $K$ is of type $F_{m}$ if and only if for every character $\chi\in\operatorname{Hom}(G,\mathbb{R})$ such that $\chi(K)=0$, we have $[\chi]\in\Sigma^{m}(G)$.
@@
Let $-\infty\leq p\leq q\leq r\leq+\infty$. If for every vertex $v\in Y^{q<\chi\leq r}$ the descending link $\operatorname{lk}_{Y^{p\leq\chi}}^{h\downarrow}(v)$ is $(k-1)$-connected, then the pair $(Y^{p\leq\chi\leq r},Y^{p\leq\chi\leq q})$ is $k$-connected. If for every vertex $v\in Y^{p\leq\chi<q}$ the ascending link $\operatorname{lk}_{Y^{\chi\leq r}}^{h\uparrow}(v)$ is $(k-1)$-connected, then the pair $(Y^{p\leq\chi\leq r},Y^{q\leq\chi\leq r})$ is $k$-connected.
@@
Let $h=(\chi,f):Y\rightarrow\mathbb{R}\times\mathbb{R}$ be a Morse function. If $Y$ is $(m-1)$-connected and for every vertex $v\in Y^{\chi<q}$ the ascending link $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is $(m-1)$-connected, then $Y^{q\leq\chi}$ is $(m-1)$-connected.
Informally, we can think of the characters on $P\mathcal{H}_{n}$ as being generated by counting the handle shifts $\rho_{1},\dots,\rho_{n-1}$, where $\rho_{i}$ shifts the $i$th end into the $n$th end by one piece. Specifically, for $11\leq i\leq n-$ 1, take $\chi_{i}(\varphi)$ to be the negative of the sum of the powers of $\rho_{i}$ appearing in $\varphi$; for $i=n$, take $\chi_{n}$ to be the sum of the powers of $\rho_{1}$ through $\rho_{n-1}$. We stress that this definition, while convenient to use, is not obviously well-defined. An equivalent definition, which is better suited to demonstrating well-definedness, is as follows (details can be found in Section 3 of [APV17]).
@@
Given a vertex, we can define the character height function $h_{\chi_{i}}([Z,\varphi])$ by adding the number of pieces of $Z$ in the $i$th end to $\chi_{i}(\varphi)$. This is well defined: consider two representatives $[Z,\varphi]=[W,\psi]$. We have that $\varphi^{-1}\psi$ takes $W$ to $Z$, and is rigid elsewhere. This composition can be further decomposed as a mapping class of some suited subsurface composed with a sequence of handle shifts. This compactly supported mapping class affects neither the distribution of the pieces of $Z$ nor the value of any character on $\varphi$, so we need consider only handle shifts. Consider $[Z,\varphi]=[\rho_{i}(Z),\varphi\circ\rho_{i}^{-1}]$: we have moved one piece from the $i$th end to the $n$th end, while increasing the amount of pushing into the $i$th end and out of the $n$th end by one. These cancel out in $h_{\chi_{i}}$ and $h_{\chi_{n}}$. All other basis characters are unchanged under these operations, and thus we have well-defined character height functions on $X_{n}$. For an arbitrary character $\chi$, we shall henceforth abuse notation by writing $\chi$ when we mean $h_{\chi}$. The domain will typically be clear.
@@ 3 The Stein–Farley complex $X_{n}$ is CAT(0)
We have a contractible cube complex on which $P\mathcal{H}_{n}$ acts nicely. In this section, we show that $X$ is in fact CAT(0). To begin, we require the following definition.
@@
In [GLU20], there is a third condition: that the 1 skeleton not contain a copy of the complete bipartite graph $K_{2,3}$. This third condition is redundant: A careful examination of the proof of Theorem 6.1 in [Che00], especially the paragraph beginning “To prove the converse”, yields that the other hypotheses imply the non-existence of $K_{2,3}$’s. Another, smaller modification: there was a condition that the 1-skeleton be a connected graph, which is automatically true in a simply connected cube complex.
@@
In order to show that $X_{n}$ is cube-complete, we shall proceed as follows: first, we show that given a vertex $[Z,\varphi]$, there is only one ascending edge per direction (see Lemma 3.5). Secondly, we show that there are no collapsed squares. This easily gives that there are no collapsed cubes of any dimension. Finally, we show that there are no empty squares, and from this induct to show that there are no empty cubes.
@@
(Uniqueness of Ascending Edges) Given a vertex $v=[Z,\varphi]$, there exists exactly one ascending edge in each direction, i.e. for i{1,,n}normal-$i\in\{1,\dots,n\}$, the collection $\{[Z\cup B_{i},\varphi]\}$, where $B_{i}$ is in the $i$th end, exhausts the ascending link of $v$.
@@
What we wish to show is that, given two representatives $[Z,\varphi]=[Z^{\prime},\varphi^{\prime}]$, whenever we ascend from $[Z^{\prime},\varphi^{\prime}]$ by adding a piece in the $i$th end, we obtain the same vertex as by adding a piece in the $i$th end to $[Z,\varphi]$. Consider the following diagram:
{tikzpicture}
Our goal is to show that, assuming $B^{\prime}$ and $B$ are both in the $i$th end, we have $[Z^{\prime}\cup B^{\prime},\varphi^{\prime}]=[Z\cup B,\varphi]$. As $(\varphi^{\prime})^{-1}\circ\varphi$ is rigid outside $Z$, it takes $B$ to some piece adjacent to $Z^{\prime}$. If this is $B^{\prime}$, then we have that $(\varphi^{\prime})^{-1}\circ\varphi$ takes $Z\cup B$ to $Z^{\prime}\cup B^{\prime}$, and is rigid elsewhere, i.e. that $[Z\cup B,\varphi]=[Z^{\prime}\cup B^{\prime},\varphi^{\prime}]$. Thus, there can be only one ascending edge per direction. ∎
Combining this with the $L_{g}$ representatives of Lemma 2.7, we see that the ascending edges over $[L_{g},\varphi]$ are of the form $[L_{g+1},\varphi\rho_{i}^{-1}]$, for $i\in\{1,\dots,n\}$. (Recall that $\rho_{n}$ is just the identity.) Next, we prove that there are no collapsed squares.
@@
For the sake of contradition, suppose there is some collapsed square, which has only two values for the heights of its vertices. Writing the lower height vertices as $[L_{g},\varphi]$ and $[L_{g},\psi]$, we see that a collapsed square must occur as in Figure 2. That the same representative can be used for both edges from one of the side vertices is justified by Lemma 3.5.
{tikzpicture}
@@
To show that the collapsed square in Figure 2 cannot occur, we assume that the middle pair of vertices are distinct, and prove that the leftmost and rightmost vertices cannot be distinct. The equality of the different forms of the middle vertices says that the transition map $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ takes $L_{g+1}$ to itself and is rigid elsewhere, for $i=1,2$. This implies that $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ has net-zero shifting in all ends, and hence is compactly supported (recall that for asymptotically rigid maps, these are the same). As the middle vertices are distinct, we know that $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$. Thus, we see that $i_{k}=j_{k}$ for $k=1,2$.
We will now show that $[L_{g},\varphi]=[L_{g},\psi]$ by considering the transition map $\varphi^{-1}\psi$ on pieces outside of $L_{g}$. The rigidity of the $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ maps outside $L_{g+1}$ immediately handles all but three pieces: $B_{i_{1}}^{1}$, $B_{i_{2}}^{1}$, and $B_{n}^{g+1}$. The last of these is again easy: the $L_{g+1}$ transition maps are rigid on the $g+2$nd piece, which they first push to the $g+1$st piece, then apply $\varphi^{-1}\psi$, then push back out by one. As this must be sent rigidly to the $g+2$nd piece, we see that $\varphi^{-1}\psi$ must be rigid on the $g+1$st piece. For $B_{i_{1}}^{1}$, consider that $\varphi^{-1}\psi$ and $\rho_{i_{2}}\varphi^{-1}\psi\rho_{i_{2}}^{-1}$ act the same on it. As the latter is rigid there, so is the former. Symmetrically, we have rigidity on $B_{i_{1}}^{1}$, and are done. Specifically, we have shown that $\varphi^{-1}\psi$ takes $L_{g}$ to itself, and is rigid elsewhere, implying that the collapsed square we began with was degenerate to begin with.
@@
We begin by observing that there can be no collapsed cubes: any such cube must contain a collapsed square. Now we show that there are no empty squares. As there are no collapsed squares, we know that the 1-skeleton of a square must have a lowest vertex, say at height $g$, two vertices at height $g+1$, and one vertex at height $g+2$. By the uniqueness (per end) of ascending edges (Lemma 3.5), we see that given a representative $[Z,\varphi]$ for the height $g$ vertex, the other vertices must be of the form $[-,\varphi]$, with the blank filled by $Z\cup B_{1}$, $Z\cup B_{2}$, and $Z\cup B_{1}\cup B_{2}$, depending on height. The only concern about the $g+2$ vertex would be that the pieces lie in the same end, but we see that this cannot occur. Thus, there are no empty squares.
Finally, we show that there are no empty cubes of any dimension. We do so inductively: suppose we already have that there are no empty $(d-1)$-cubes. Consider the $1$-skeleton of an $d$-cube $C$. Let the bottom vertex of $C$ be $[Z,\varphi]$, at height $g$. Each of the vertices of $C$ adjacent to $[Z,\varphi]$ is of the form $[Z\cup B_{i},\varphi]$, where $B_{i}$ is in the $i$th end; each $i$ appears at most once by Lemma 3.5. Also, by the inductive hypothesis, we have that $[Z,\varphi]$ connects to each of the vertices at height $g+d-1$ via a $(d-1)$-cube, so that these vertices can be written as $[Z\cup B_{i_{1}}\cup\cdots B_{i_{j-1}}\cup\widehat{B}_{i_{j}}\cup B_{i_{j+1}}\cup\cdots\cup B_{i_{d}},\varphi]$, where the hat indicates omission. Choose two such vertices, and consider their common lower vertex (the one missing the two pieces missing in either of the chosen vertices). Applying the doctrine of no empty squares to these 3 vertices and the apex yields that the apex can be written as $[Z\cup B_{i_{1}}\cup\cdots\cup B_{i_{d}},\varphi]$, and thus the cube is filled.
@@
The argument above that ascending edges are unique per direction emphatically does not hold for descending edges. In fact, for any vertex of height at least 1, there are infinitely many descending edges per direction in which it can have a descending edge (given $[Z,\varphi]$, choose $\psi\in\mathcal{PH}_{n}$ such that $\varphi^{-1}\psi|_{Z}$ is a homeomorphism, with some nontrivial behavior in the piece to be removed, and such that $\varphi^{-1}\psi$ is rigid outside $Z$). It is helpful to keep in mind the following: when ascending, one can always choose coherent representatives from a minimal vertex; when descending, nothing is guaranteed.
@@
{tikzcd}
The difficulty is in the remaining case, where the common vertex is maximal in the cube. We begin with the diagram in Fig. 3. Our goal is to obtain a representative $[L_{g+2},\varphi\circ\rho_{k}]$ for the height $g+2$ vertex currently represented with both $\varphi^{\prime}$ and $\varphi^{\prime\prime}$. This will imply the existence of a vertex $[L_{g},\varphi\circ\rho_{i}\rho_{j}\rho_{k}]$, which realizes this 3-wheel in a 3-cube. To begin, we have the following lemma that justifies part of the diagram.
@@
Whenever $[L_{g},\varphi\circ\rho_{i}]=[L_{g},\varphi^{\prime}\circ\rho_{j}]$ and $[L_{g+1},\varphi]=[L_{g+1},\varphi^{\prime}]$, we have that $i=j$.

### v2

@@ 2.3 Characters and $\Sigma$-Invariants
Recall that a group is of type Fm if it admits a proper cocompact action on an (m1)(m-1)-connected CW-complex. Assume $Y$ is a CW complex and $h:Y\rightarrow\mathbb{R}$ is continuous. We call the corresponding filtration $\left(Y^{t\leq h}\right)_{t\in\mathbb{R}}$ on $Y$ essentially (m1m-1)-connected if for any $t\in\mathbb{R}$, there exists $s\leq t$ such that inclusion $Y^{t\leq h}\hookrightarrow Y^{s\leq h}$ induces the trivial map in $\pi_{k}$ for $k\leq m-1$.
For any group $G$, we define a character of $G$ to be a homomorphism from $G$ to the additive group $\mathbb{R}$. As characters factor through $G^{ab}$ (the abelianization of $G$), the space of characters is a vector space with the same dimension as the rank of $G^{ab}$. Excluding the trivial character and modding out by positive scaling, we obtain the Character Sphere $\Sigma(G)$. The $\Sigma$-invariants (also called BNS or BNSR invariants, for Bieri-Neumann-Strebel(-Renz)) are a filtration of the character sphere into subspaces
@@
Let $G$ be a group of type $F_{m}$, and let $Y$ be an (m1)(m-1)-connected CW complex on which $G$ acts cocompactly. Suppose that the stabilizer of any $k$-cell is of type $F_{m-k}$ and is contained in the kernel of every character of GG. [^2] That the cell-stabilizers are contained in the kernels is the modification. For each non-trivial $\chi\in\operatorname{Hom}(G,\mathbb{R})$, there is a character height function $h_{\chi}:Y\rightarrow\mathbb{R}$, a continuous map satisfying $h_{\chi}(gy)=\chi(g)+h_{\chi}(y)$ for all $y\in Y$ and $g\in G$. Then $\Sigma^{m}(G)$ is the set of those $\chi]$ such that the filtration $(Y^{t\leq h_{\chi}})_{t\in\mathbb{R}}$ is essentially (m1)(m-1)-connected.
Of what use are these invariants, one might ask? A standard result is the following classification of when finiteness properties of $G$ are preserved to subgroups containing the commutator subgroup.
@@
Let $G$ be a group of type $F_{m}$, and let $K$ be a subgroup so that $G,G]\leq K\leq G$. Then $K$ is of type $F_{m}$ if and only if for every character $\chi\in\operatorname{Hom}(G,\mathbb{R})$ such that $\chi(K)=0$, we have $\chi]\in\Sigma^{m}(G)$.
@@
Let $\infty\leq p\leq q\leq r\leq+\infty$. If for every vertex $v\in Y^{q<\chi\leq r}$ the descending link $\operatorname{lk}_{Y^{p\leq\chi}}^{h\downarrow}(v)$ is (k1)(k-1)-connected, then the pair $(Y^{p\leq\chi\leq r},Y^{p\leq\chi\leq q})$ is $k$-connected. If for every vertex $v\in Y^{p\leq\chi<q}$ the ascending link $\operatorname{lk}_{Y^{\chi\leq r}}^{h\uparrow}(v)$ is (k1)(k-1)-connected, then the pair $(Y^{p\leq\chi\leq r},Y^{q\leq\chi\leq r})$ is $k$-connected.
@@
Let $h=(\chi,f):Y\rightarrow\mathbb{R}\times\mathbb{R}$ be a Morse function. If $Y$ is (m1)(m-1)-connected and for every vertex $v\in Y^{\chi<q}$ the ascending link $\operatorname{lk}_{Y}^{h\uparrow}(v)$ is (m1)(m-1)-connected, then $Y^{q\leq\chi}$ is (m1)(m-1)-connected.
Informally, we can think of the characters on $P\mathcal{H}_{n}$ as being generated by counting the handle shifts $\rho_{1},\dots,\rho_{n-1}$, where $\rho_{i}$ shifts the $i$th end into the $n$th end by one piece. Specifically, for $1\leq i\leq n-1$, take $\chi_{i}(\varphi)$ to be the negative of the sum of the powers of $\rho_{i}$ appearing in $\varphi$; for $i=n$, take $\chi_{n}$ to be the sum of the powers of $\rho_{1}$ through $\rho_{n-1}$. We stress that this definition, while convenient to use, is not obviously well-defined. An equivalent definition, which is better suited to demonstrating well-definedness, is as follows (details can be found in Section 3 of [APV17]).
@@
Given a vertex, we can define the character height function $h_{\chi_{i}}([Z,\varphi])$ by adding the number of pieces of $Z$ in the $i$th end to $\chi_{i}(\varphi)$. This is well defined: consider two representatives $Z,\varphi]=[W,\psi]$. We have that $\varphi^{-1}\psi$ takes $W$ to $Z$, and is rigid elsewhere. This composition can be further decomposed as a mapping class of some suited subsurface composed with a sequence of handle shifts. This compactly supported mapping class affects neither the distribution of the pieces of $Z$ nor the value of any character on $\varphi$, so we need consider only handle shifts. Consider $Z,\varphi]=[\rho_{i}(Z),\varphi\circ\rho_{i}^{-1}]$: we have moved one piece from the $i$th end to the $n$th end, while increasing the amount of pushing into the $i$th end and out of the $n$th end by one. These cancel out in $h_{\chi_{i}}$ and $h_{\chi_{n}}$. All other basis characters are unchanged under these operations, and thus we have well-defined character height functions on $X_{n}$. For an arbitrary character $\chi$, we shall henceforth abuse notation by writing $\chi$ when we mean $h_{\chi}$. The domain will typically be clear.
@@ 3 The Stein–Farley complex $X_{n}$ is CAT(0)
We have a contractible cube complex on which $P\mathcal{H}_{n}$ acts nicely. In this section, we show that $X_{n}$ is in fact CAT(0). To begin, we require the following definition.
@@
In [GLU20], there is a third condition: that the 1 skeleton not contain a copy of the complete bipartite graph $K_{2,3}$. This third condition is redundant: A careful examination of the proof of Theorem 6.1 in [Che00], especially the paragraph beginning “To prove the converse”, yields that the other hypotheses imply the non-existence of $K_{2,3}$’s: A $K_{2,3}$ is a 3-wheel where all the vertices opposite the common vertex are identified, which cannot be the case in an embedded 3-cube. Another, smaller modification: there was a condition that the 1-skeleton be a connected graph, which is automatically true in a simply connected cube complex.
@@
In order to show that $X_{n}$ is cube-complete, we shall proceed as follows: first, we show that given a vertex $Z,\varphi]$, there is only one ascending edge per direction (see Lemma 3.5). Secondly, we show that there are no collapsed squares. This easily gives that there are no collapsed cubes of any dimension. Finally, we show that there are no empty squares, and from this induct to show that there are no empty cubes.
@@
(Uniqueness of Ascending Edges) Given a vertex $v=[Z,\varphi]$, there exists exactly one ascending edge in each direction, i.e. for $i\in\{1,\dots,n\}$, the collection $\{[Z\cup B_{i},\varphi]\}$, where $B_{i}$ is the piece in the $i$th end adjacent to $Z$, exhausts the vertices of the ascending link of $v$.
@@
What we wish to show is that, given two representatives $Z,\varphi]=[Z^{\prime},\varphi^{\prime}]$, whenever we ascend from $Z^{\prime},\varphi^{\prime}]$ by adding a piece in the $i$th end, we obtain the same vertex as by adding a piece in the $i$th end to $Z,\varphi]$. Consider the following diagram:
$Z,\varphi][$$Z^{\prime},\varphi^{\prime}][$$Z\cup B,\varphi][$$Z^{\prime}\cup B^{\prime},\varphi^{\prime}]$?
Our goal is to show that, assuming $B^{\prime}$ and $B$ are both in the $i$th end, we have $Z^{\prime}\cup B^{\prime},\varphi^{\prime}]=[Z\cup B,\varphi]$. As $(\varphi^{\prime})^{-1}\circ\varphi$ is rigid outside $Z$, it takes $B$ to some piece adjacent to $Z^{\prime}$. As $(\varphi^{\prime})^{-1}\circ\varphi\in P\mathcal{H}_{n}$, this implies that $(\varphi^{\prime})^{-1}\circ\varphi(B)$ must still be in the $i$th end (elements of $P\mathcal{H}_{n}$ cannot rigidly move pieces from one end into another), and hence is $B^{\prime}$. We then have that $(\varphi^{\prime})^{-1}\circ\varphi$ takes $Z\cup B$ to $Z^{\prime}\cup B^{\prime}$, and is rigid elsewhere, i.e. that $Z\cup B,\varphi]=[Z^{\prime}\cup B^{\prime},\varphi^{\prime}]$. Thus, there can be only one ascending edge per direction. ∎
Combining this with the $L_{g}$ representatives of Lemma 2.7, we see that the ascending edges over $L_{g},\varphi]$ are of the form $L_{g+1},\varphi\rho_{i}^{-1}]$, for $i\in\{1,\dots,n\}$. (Recall that $\rho_{n}$ is just the identity.) Next, we prove that there are no collapsed squares.
@@
For the sake of contradiction, suppose there is some collapsed square, which has only two values for the heights of its vertices. Writing the lower height vertices as $L_{g},\varphi]$ and $L_{g},\psi]$, we see that a collapsed square must occur as in Figure 2. That the same representative can be used for both edges from one of the side vertices is justified by Lemma 3.5.
$L_{g},\varphi][$$L_{g+1},\varphi\rho_{i_{1}}^{-1}][$$L_{g+1},\psi\rho_{j_{1}}^{-1}][$$L_{g+1},\varphi\rho_{i_{2}}^{-1}][$$L_{g+1},\psi\rho_{j_{2}}^{-1}][$$L_{g},\psi]$
@@
To show that the collapsed square in Figure 2 cannot occur, we assume that the middle pair of vertices are distinct, and prove that the leftmost and rightmost vertices cannot be distinct. The equality of the different forms of the middle vertices says that the transition map $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ takes $L_{g+1}$ to itself and is rigid elsewhere, for $k=1,2$. This implies that $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ has net-zero shifting in all ends, and hence is compactly supported (recall that for asymptotically rigid maps, these are the same). As the middle vertices are distinct, we know that $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$. Thus, we see that $i_{k}=j_{k}$ for $k=1,2$.
We will now show that $L_{g},\varphi]=[L_{g},\psi]$ by considering the transition map $\varphi^{-1}\psi$ on pieces outside of $L_{g}$. The rigidity of the $\rho_{i_{k}}\varphi^{-1}\psi\rho_{j_{k}}^{-1}$ maps outside $L_{g+1}$ immediately handles all but three pieces: $B_{i_{1}}^{1}$, $B_{i_{2}}^{1}$, and $B_{n}^{g+1}$. The last of these is again easy: the $L_{g+1}$ transition maps are rigid on the g+2g+2nd piece, which they first push to the g+1g+1st piece, then apply $\varphi^{-1}\psi$, then push back out by one. As this must be sent rigidly to the g+2g+2nd piece, we see that $\varphi^{-1}\psi$ must be rigid on the g+1g+1st piece. For $B_{i_{1}}^{1}$, consider that $\varphi^{-1}\psi$ and $\rho_{i_{2}}\varphi^{-1}\psi\rho_{i_{2}}^{-1}$ act the same on it. As the latter is rigid there, so is the former. Symmetrically, we have rigidity on $B_{i_{2}}^{1}$, and are done. Specifically, we have shown that $\varphi^{-1}\psi$ takes $L_{g}$ to itself, and is rigid elsewhere, implying that the collapsed square we began with was degenerate to begin with.
@@
We begin by observing that there can be no collapsed cubes: any such cube must contain a collapsed square. Now we show that there are no empty squares. As there are no collapsed squares, we know that the 1-skeleton of a square must have a lowest vertex, say at height $g$, two vertices at height $g+1$, and one vertex at height $g+2$. By the uniqueness (per end) of ascending edges (Lemma 3.5), we see that given a representative $Z,\varphi]$ for the height $g$ vertex, the other vertices must be of the form $\varphi]$, with the blank filled by $Z\cup B_{1}$, $Z\cup B_{2}$, and $Z\cup B_{1}\cup B_{2}$, depending on height. The only concern about the $g+2$ vertex would be that the pieces lie in the same end, but we see that this cannot occur. Thus, there are no empty squares.
Finally, we show that there are no empty cubes of any dimension. We do so inductively: suppose we already have that there are no empty (d1)(d-1)-cubes. Consider the $1$-skeleton of a $d$-cube $C$. Let the bottom vertex of $C$ be $Z,\varphi]$, at height $g$. Each of the vertices of $C$ adjacent to $Z,\varphi]$ is of the form $Z\cup B_{i},\varphi]$, where $B_{i}$ is in the $i$th end; each $i$ appears at most once by Lemma 3.5. Also, by the inductive hypothesis, we have that $Z,\varphi]$ connects to each of the vertices at height g+d1g+d-1 via a (d1)(d-1)-cube, so that these vertices can be written as $Z\cup B_{i_{1}}\cup\cdots B_{i_{j-1}}\cup\widehat{B}_{i_{j}}\cup B_{i_{j+1}}\cup\cdots\cup B_{i_{d}},\varphi]$, where the hat indicates omission. Choose two such vertices, and consider their common lower vertex (the one missing the two pieces missing in either of the chosen vertices). Applying the doctrine of no empty squares to these 3 vertices and the apex yields that the apex can be written as $Z\cup B_{i_{1}}\cup\cdots\cup B_{i_{d}},\varphi]$, and thus the cube is filled.
@@
The argument above that ascending edges are unique per direction emphatically does not hold for descending edges. In fact, for any vertex of height at least 1, there are infinitely many descending edges per direction in which it can have a descending edge (given $Z,\varphi]$, choose $\psi\in\mathcal{PH}_{n}$ such that $\varphi^{-1}\psi|_{Z}$ is a homeomorphism, with some nontrivial behavior in the piece to be removed, and such that $\varphi^{-1}\psi$ is rigid outside $Z$). It is helpful to keep in mind the following: when ascending, one can always choose coherent representatives from a minimal vertex; when descending, nothing is guaranteed.
@@
$\cdot$$\cdot$$\cdot$$\cdot$$\cdot$$\cdot$$\cdot$$\cdot$}
The difficulty is in the remaining case, where the common vertex is maximal in the cube. We begin with the diagram in Fig. 3. Our goal is to obtain a representative $L_{g+2},\varphi\circ\rho_{k}]$ for the height $g+2$ vertex currently represented with both $\varphi^{\prime}$ and $\varphi^{\prime\prime}$. This will imply the existence of a vertex $L_{g},\varphi\circ\rho_{i}\rho_{j}\rho_{k}]$, which realizes this 3-wheel in a 3-cube. To begin, we have the following lemma that justifies part of the diagram.
@@
Whenever $L_{g},\varphi\circ\rho_{i}]=[L_{g},\varphi^{\prime}\circ\rho_{j}]$ and $L_{g+1},\varphi]=[L_{g+1},\varphi^{\prime}]$, we have that $i=j$.